OpenStudy (anonymous):

Prove that every prime number is the leg of exactly one right triangle with integer sides. ~I've noticed that pythagorean triples usually contain atleast one prime number, but I am not exactly sure how you would go about proving that.

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

found it?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

oh. Ok one second

myininaya (myininaya):

Euclid's formula[1] is a fundamental formula for generating Pythagorean triples given an arbitrary pair of positive integers m and n with m > n. The formula states that the integers maybe we can use this

myininaya (myininaya):

oops it didn't copy the formulas

myininaya (myininaya):

a=m^2-n^2 b=2mn c=m^2+n^2

OpenStudy (anonymous):

So when using this formula would we be subbing in prime numbers for m and n?

myininaya (myininaya):

a=k(m^2-n^2) b=k(2mn) c=k(m^2+n^2) k is a positive integer no m and n are integers (can be anything)

OpenStudy (anonymous):

Okay~ i think I'm following you..

myininaya (myininaya):

like i said i'm not sure if we will need this our not im still thinking

myininaya (myininaya):

http://mathforum.org/library/drmath/view/67289.html this might help